OpenStudy (kimberly101):

A landscape gardener wishes to spread fertilizer on a semicircular garden bed which has a diameter of 4.7m. How much fertilizer is required, if the fertilizer is applied at the rate of 20g per square metre.

OpenStudy (mysesshou):

circle area = \[pi*r^2\] semicircle area = \[0.5*pi*r^2\] d=4.7m r=0.5*d = 2.35m semicircle area = \[0.5*pi*(2.35m)^2 = 8.674m^2\] \[(8.674m^2)(20g/m^2) = 173.48g\]

OpenStudy (kimberly101):

thanks!!!